OpenStudy (anonymous):

prime Factors of x3-1

OpenStudy (anonymous):

\[x^3-1=(x-1)(x^2+x+1)\]

OpenStudy (anonymous):

Can you explain how you get your answer please :]]

OpenStudy (lgbasallote):

difference of two cubes a^3 = b^3 = (a-b)(a^2 + ab + b^2)

OpenStudy (anonymous):

how about x3-8 ?